OpenStudy (anonymous):

Just need the answer plz!!! Will medal! How many solutions does the equation have? A. one solution B. infinite solutions C. no solution

OpenStudy (kamibug):

What's the equation?

OpenStudy (anonymous):

Oh sorry It's 3(4x-2)=2(6x-3)

OpenStudy (anonymous):

12x-6=12x-6 = 0 so

OpenStudy (anonymous):

Ok thx just checking my answers((:

OpenStudy (anonymous):

You welcome

OpenStudy (anonymous):

So it woud be A or C?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

Meant A

OpenStudy (anonymous):

Thx